OpenStudy (anonymous):

The storage shed shape of a half-cylinder of radius r and length l. The shed is filled with sawdust whose density (mass/unit volume) at any point is proportional to the distance of that point from the floor. The constant of proportionality is k. Calculate the total mass of sawdust in the shed.

OpenStudy (koikkara):

So, here as per the question, The shed is filled with sawdust whose density at any point is proportional to the distance of that point from the floor. the constant of proportionality is k. calculate the total mass of sawdust in the shed. Volume of the Shed = (πr² /2 )l Given: Density = kd ,d = distance from the floor Density = Mass (M) / Volume kd = M / (πr² /2)l M = π r² l / 2 (kd) >>>>> M = π l r² kd/2 << Ans....... well...i think this is it....lol

OpenStudy (anonymous):

its a half cylinder so volume is πr^2l / 2

OpenStudy (anonymous):

my problem is that i dont know how to find "d"

OpenStudy (anonymous):

mass = πr^2l/2 *k*d(distance from floor)

OpenStudy (anonymous):

I need to figure out what the distance from the floor is

OpenStudy (koikkara):

density (mass/unit volume)

OpenStudy (anonymous):

and density = k * (distance from floor)

OpenStudy (koikkara):

well, it may help than i could.... http://mathforum.org/kb/message.jspa?messageID=688054

OpenStudy (koikkara):

hope that defines evrythin....

OpenStudy (anonymous):

doesn't explain how to find "d" :(

OpenStudy (koikkara):

well, one user is there who can explain them...his name is "In Your Head"....he can sort out, as i rembr only some formulas and not in detail...well, i passed out 12th...4yrs ago...

OpenStudy (koikkara):

@Ldaniel ...ask...user named..."In Your Head"...

OpenStudy (anonymous):

@inyourhead

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

there might be some integration needed.

OpenStudy (anonymous):

ok this is what i have so far \[Volume = \frac{\pi*r^{2}*L}{2}\] \[Density = kd\] \[Density = \frac{mass}{volume}\] \[kd = \frac{mass}{\frac{\pi*r^{2}*L}{2}}\] \[mass = \frac{2*k*d}{\pi*r^{2}*L}\]

OpenStudy (anonymous):

still you didn't find d

OpenStudy (anonymous):

true. however, I believe that some calculus is needed as Koikkara's link shows.

OpenStudy (anonymous):

this may help http://mathhelpforum.com/calculus/130945-finding-mass-density-volume.html